Designer Jackets, Inc. produces blazers and has the following data available on these products.

Sales price per unit

$ 280

Variable cost per unit

120

Fixed costs per month

16,000

Tax rate

20 percent

(1) How many units must be sold to earn a monthly profit of $150,000 after taxes?
(2) How many sales dollars are required to earn a monthly profit of $150,000 after taxes?
Answer #1

Answer :

1) 1,272 (or) 1271.875 Units

2) $ 356,125

Contribution Margin = Selling Price - Vara ible cost = 280-120 = 160

CM ratio = 160/280 = 57.14%

Units Required to Get the $150,000 After Tax Profit

= (Fixed cost +( ( Profit )/ (1-Tax) ) / Contribution MArgin = ((16,000+(150000/(1-20%)) / 160 = 1271.875 Units (Answer)

Dollars Required = 1,271.875*280 = $ 356,125 (Answer)